OpenStudy (anonymous):

Easton simplified the following expression: (x2y6z5)(x4y5z3) If he writes his answer in the form of xaybzc, what is the value of b, the exponent on y?

OpenStudy (helder_edwin):

just add the exponents of y you already have remember this \[ a^na^m=a^{n+m} \]

OpenStudy (anonymous):

\[\Large (x^{2}y^{6}z^{5})(x^{4}y^{5}z^{3}) = (x^{2}x^{4})(y^{6}y^{5})(z^{5}z^{3})\] Use the wquation @helder_edwin posted.

OpenStudy (anonymous):

\[a^m a^n=a^{m+n}\] Figure it out.

OpenStudy (anonymous):

.... x^2y^5z^3(yz^2+x^2) ?

OpenStudy (anonymous):

No idea how you got that. You add the EXPONENTS, not the variables.

OpenStudy (anonymous):

Example: \[\LARGE T^{4}T^{7} = T^{4+7} = T^{11}\]

OpenStudy (anonymous):

okay i get it now .

OpenStudy (anonymous):

so the answer is?......